# Understanding Dynamic Human-Robot Proxemics in the Case of Four-Legged Canine-Inspired Robots

### v4

@@ I Introduction
In this paper, we present a proxemic user study ($N$=32) involving a canine robot in a motion capture environment. The study involves a dynamic scenario where participants pass by the moving canine robot while engaged in an unrelated task, with their positions recorded in real-time by motion capture cameras for subsequent analysis. The objective of this experiment is to explore the impact of different interaction postures on participants’ proxemic behavior, contributing to the broader understanding of human-robot interaction in dynamic environments.
@@ II-A1 Human Proxemics
The study of proxemic relationships, which explores the spatial behavior between individuals, has been a focus of research for many years [11, 12]. Hall [13] introduced the concept of proxemics, categorizing personal space into four distinct zones: intimate (6–18 inches) for close relationships, personal (1.5–4 feet) for family and close friends, social (4–12 feet) for acquaintances, and public (12–25 feet) for public settings. These zones help communicate varying levels of intimacy and relational distance during interactions.
Aiello [14] later expanded on this concept, suggesting that proxemic behavior serves two primary purposes: 1) to mitigate strong emotional responses triggered by close physical proximity [15], and 2) to maintain sufficient space to manage potential threats [16, 10]. Proxemic distances are not static; they change depending on the nature of the interaction and the entities involved. The interpersonal distances people naturally maintain reflect their attitudes toward others or objects in their environment, which in turn influences their behavior.
@@ II-A2 Human-Robot Proxemics
As robots become more integrated into daily life, particularly in private spaces, understanding and adhering to human proxemic norms becomes crucial. This is especially true for robots used in home services [20, 21]. Mumm et al. [4] argue that robots must be designed to follow societal norms regarding physical and psychological distancing to interact effectively with humans. A key aspect of this is the manipulation of the robot gaze, which plays a critical role in psychological proxemics. Additionally, the static orientation, distance, and signaling of robots are significant factors that influence how humans perceive and interact with them [8]. Eresha et al. [22] found that unwanted gaze during human-robot interactions can cause discomfort or distract the human participant, highlighting the importance of gaze in designing socially acceptable robots.
@@ II-B Movement Capture in Proxemics
Various methods have been developed for studying indoor proxemics, including RFID [27], virtual reality (VR) [10], and wearable devices [28]. Although these technologies offer convenience and mobility, none match the precision of motion capture systems, which are widely used in the creation of state-of-the-art virtual reality experiences and 3D films due to their high precision, quick response, and high sampling rates [29, 30, 31]. Research using motion capture systems in proxemics studies leverages the system’s capability to analyze participants’ movements at all key body points [32], extracting features relevant to emotion or perception. For instance, Mead et al. [11] used a Hidden Markov Model to recognize participants’ behaviors based on motion data collected from a Microsoft Kinect camera. Jakobsen et al. [33] focused on position, velocity, and orientation to find efficient methods for information visualization in the context of proxemics.
@@ III-C2 Task Execution
(a) Refer to caption (b) Refer to caption
Figure 5: (a): Trajectories of a single participant. The orange fixed trajectory for Spot, other colored trajectories for humans, and the time-synchronized minimum distances are labeled as dotted lines in the plot. (b): Distance in bar plot corresponding to trajectories in (a).
@@ III-D Data Collection
We recruited 32 participants (17 males, 15 females), with ages ranging from 19 to 37 years ($M=26,SD=3.57$). Their experience with robots varied. This study was approved by our ethics committee. Each participant underwent 2$\times$4 = 8 different conditions, with the order of conditions counterbalanced using an 8$\times$8 Latin square. Participants were not informed of the sequence of cases.
@@ IV-B Hypotheses and Results
Based on the results we observed from the above distances, we made several assumptions and performed a two-way $2*4$ Analysis of variance (ANOVA) with two within-participant factors: the robot gaze factor of two levels (with/without), and the robot orientation of four levels (stationary, forward, sideways and backward). The dependent variable is the time-synchronized minimum participant-robot distance for each set of conditions. We verified the data on the normality assumption (Shapiro-Wilk test, $p>0.05$) and the homogeneity of variances (Levene’s test, $p>0.05$). The ANOVA test revealed a significant main effect of the robot orientation on distance ($F(3,256)=7.53,p=0.00008,partial~{}\eta^{2}=0.091$). Gaze also showed a significant main effect ($F(1,256)=21.89,p=0.000005,partial~{}\eta^{2}=0.089$). However, the interaction effect between orientation and gaze was not significant, ($F(3,256)=0.69,p=0.56,partial~{}\eta^{2}=0.009$).
@@
Our results, however, offer a different perspective. The experienced group ($N=12,M=1.24,SD=0.38$) maintained a slightly longer mean distance than the inexperienced group ($N=20,M=1.16,SD=0.36$), as shown in Fig. 6(c). Interestingly, this observation contradicts the hypothesis, suggesting that more experienced participants might tend to be more cautious toward the quadruped robot, possibly due to their deeper understanding of robotic behaviors and potential risks. In addition, we performed a t-test to see the effect of experience on personal distances, and the result indicated that this difference was not significant ($t$-$statistic=-1.56,p=0.12$). This lack of significance may be attributed to the small sample size of experienced participants ($N=12$), which limits the power of the statistical analysis. While the results do not provide strong evidence to support H4, they do highlight an interesting trend that warrants further investigation with a larger sample size. The finding that experienced participants may maintain greater distances could be influenced by their knowledge and understanding of the potential unpredictability or operational nuances of robots like the Spot.

### v5

@@ I Introduction
In this paper, we present a proxemic user study (NN=32) involving a canine robot in a motion capture environment. The study involves a dynamic scenario where participants pass by the moving canine robot while engaged in an unrelated task, with their positions recorded in real-time by motion capture cameras for subsequent analysis. The objective of this experiment is to explore the impact of different interaction postures on participants’ proxemic behavior, contributing to the broader understanding of human-robot interaction in dynamic environments.
@@ II-A1 Human Proxemics
The study of proxemic relationships, which explores the spatial behavior between individuals, has been a focus of research for many years [11, 12]. The authors in [13] introduced the concept of proxemics, categorizing personal space into four distinct zones: intimate (6–18 inches) for close relationships, personal (1.5–4 feet) for family and close friends, social (4–12 feet) for acquaintances, and public (12–25 feet) for public settings. These zones help communicate varying levels of intimacy and relational distance during interactions.
The authors in [14] later expanded on this concept, suggesting that proxemic behavior serves two primary purposes: 1) to mitigate strong emotional responses triggered by close physical proximity [15], and 2) to maintain sufficient space to manage potential threats [16, 10]. Proxemic distances are not static; they change depending on the nature of the interaction and the entities involved. The interpersonal distances people naturally maintain reflect their attitudes toward others or objects in their environment, which in turn influences their behavior.
@@ II-A2 Human-Robot Proxemics
As robots become more integrated into daily life, particularly in private spaces, understanding and adhering to human proxemic norms becomes crucial. This is especially true for robots used in home services [20, 21]. The authors in [4] argue that robots must be designed to follow societal norms regarding physical and psychological distancing to interact effectively with humans. A key aspect of this is the manipulation of the robot gaze, which plays a critical role in psychological proxemics. Additionally, the static orientation, distance, and signaling of robots are significant factors that influence how humans perceive and interact with them [8]. The authors in [22] found that unwanted gaze during human-robot interactions can cause discomfort or distract the human participant, highlighting the importance of gaze in designing socially acceptable robots.
@@ II-B Movement Capture in Proxemics
Various methods have been developed for studying indoor proxemics, including RFID [27], virtual reality (VR) [10], and wearable devices [28]. Although these technologies offer convenience and mobility, none match the precision of motion capture systems, which are widely used in the creation of state-of-the-art virtual reality experiences and 3D films due to their high precision, quick response, and high sampling rates [29, 30, 31]. Research using motion capture systems in proxemics studies leverages the system’s capability to analyze participants’ movements at all key body points [32], extracting features relevant to emotion or perception. For instance, the authors in [11] used a Hidden Markov Model to recognize participants’ behaviors based on motion data collected from a Microsoft Kinect camera. The authors in [33] focused on position, velocity, and orientation to find efficient methods for information visualization in the context of proxemics.
@@ III-C2 Task Execution
(a) Refer to caption
Figure 5: (a): Trajectories of a single participant. The orange fixed trajectory for Spot, other colored trajectories for humans, and the time-synchronized minimum distances are labeled as dotted lines in the plot.
@@ III-D Data Collection
[…]
@@ IV-B Hypotheses and Results
Based on the results we observed from the above distances, we made several assumptions and performed a two-way 242*4 Analysis of variance (ANOVA) with two within-participant factors: the robot gaze factor of two levels (with/without), and the robot orientation of four levels (stationary, forward, sideways and backward). The dependent variable is the time-synchronized minimum participant-robot distance for each set of conditions. We verified the data on the normality assumption (Shapiro-Wilk test, $p>0.05$) and the homogeneity of variances (Levene’s test, $p>0.05$). The ANOVA test revealed a significant main effect of the robot orientation on distance ($F(3,256)=7.53,p=0.00008,partial~\eta^{2}=0.091$). Gaze also showed a significant main effect ($F(1,256)=21.89,p=0.000005,partial~\eta^{2}=0.089$). However, the interaction effect between orientation and gaze was not significant, ($F(3,256)=0.69,p=0.56,partial~\eta^{2}=0.009$).
@@
Our results, however, offer a different perspective. The experienced group ($N=12,M=1.24,SD=0.38$) maintained a slightly longer mean distance than the inexperienced group ($N=20,M=1.16,SD=0.36$), as shown in Fig. 6(c). Interestingly, this observation contradicts the hypothesis, suggesting that more experienced participants might tend to be more cautious toward the quadruped robot, possibly due to their deeper understanding of robotic behaviors and potential risks. In addition, we performed a t-test to see the effect of experience on personal distances, and the result indicated that this difference was not significant ($t$-statistic=1.56,p=0.12statistic=-1.56,p=0.12). This lack of significance may be attributed to the small sample size of experienced participants ($N=12$), which limits the power of the statistical analysis. While the results do not provide strong evidence to support H4, they do highlight an interesting trend that warrants further investigation with a larger sample size. The finding that experienced participants may maintain greater distances could be influenced by their knowledge and understanding of the potential unpredictability or operational nuances of robots like the Spot.
@@ VI Conclusion and Future Work
VII Acknowledgement
This work was supported by the UKRI Centre for Doctoral Training in Socially Intelligent Artificial Agents, Grant Number EP/S02266X/1.